# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 2.3 Point Identification
Proposition 1 characterizes the sharp identified set for $\theta_{0}$ by only imposing Assumption 1. This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$ under support conditions on the exogenous covariate $Z_{it}$. We focus on the scenario where the endogenous covariate $X_{it}$ is discrete $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$ and there are only two periods $T=2$.
@@
$\mathcal{Z}_{1}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=0,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=x\mid z_{12})\Big{\}},$
$\mathcal{Z}_{2}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=1,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=x\mid z_{12})\Big{\}}.$
@@
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
@@
Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3}\cap\mathcal{Z}_{2}\neq\emptyset$ or $\mathcal{Z}^{j}_{4}\cap\mathcal{Z}_{1}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
When the endogenous variable $X_{it}$ is a scalar, e.g., the lagged dependent variable $X_{it}=Y_{i,t-1}$, then the definition of the two sets $\mathcal{Z}^{j}_{3},\mathcal{Z}^{j}_{4}$ can be simplified as there existing $x_{1}\neq x_{2}$ such that the corresponding choice probability is observed. Besides the sign of $\gamma_{0}$, the identification results can also bound the value of $\gamma_{0}$ from variation in the exogenous covariates.
When $X_{it}$ is multi-dimensional such as including two lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2})$ with $\gamma_{0}=(\gamma_{0}^{1},\gamma_{0}^{2})$, then $\gamma_{0}^{1}$ is identified when the required choice probability in the two sets $\mathcal{Z}^{1}_{3},\mathcal{Z}^{1}_{4}$ are observed for $(Y_{i,1},Y_{i,0})=(1,1),(Y_{i,2},Y_{i,1})=(0,1)$ or $(Y_{i,1},Y_{i,0})=(0,0),(Y_{i,2},Y_{i,1})=(1,0)$. We provide general sufficient conditions to identify the sign of $\gamma_{0}$, which may be stronger than necessary and can be relaxed in certain scenarios. For example, when we know that $\gamma_{0}^{1}+\gamma_{0}^{2}>0$ while $\gamma_{0}^{1}<0$, we can infer that $\gamma_{0}^{2}>0$ without requiring additional assumptions on the two sets $\mathcal{Z}^{2}_{3},\mathcal{Z}^{2}_{4}$.
3 Ordered Response Model
3.1 Model and Identification
Consider a setting where the outcome variable $Y_{it}$ is discrete and takes $J$ values: $Y_{it}\in\{y_{1},..,y_{J}\}$. For example, $Y_{it}$ can be different categories of income, health outcomes, or educational attainment. We study the following panel ordered response model:
$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it},$ (8)
@@ 3.1 Model and Identification
where $Y_{it}^{*}$ denotes the latent dependent variable, $b_{1}=-\infty,b_{J+1}=+\infty$, and the remaining threshold parameters $b_{j}$ (where $b_{j}\leq b_{j+1}$) can be either known or unknown for $12\leq j\leq J-$ 1. The binary choice model in (1) is nested with $J=2$ and $b_{2}=0$. Similar to model (1), $X_{it}$ is the potentially endogenous covariate that can be arbitrarily dependent on $(\alpha_{i},\epsilon_{it})$, such as the lagged dependent variable $Y_{i,t-1}$, while $Z_{it}$ is the exogenous covariate that satisfies Assumption 1:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
We now show how our key identification strategy based on partial stationarity can also be exploit to partially identify $\theta_{0}$. Similar to Section 2, we still seek to establish both lower and upper bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, where $v_{it}:=\epsilon_{it}+\alpha_{i}$. The distinction is that in the ordered choice setting we can obtain $J$ different (upper and lower) bounds, which can then be aggregated over to form tighter bounds as shown below.
1. Lower bound. By the monotonicity of a distribution, we first derive a lower bound for the distribution $v_{t}\mid w_{st}$ conditional on all covariates $W_{ist}=w_{st}$, given as
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=\Pr(v_{t}\leq a\mid w_{st})$
$\geq\Pr(v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
where the above lower bound holds for any choice $j\leq J$. Since the interval $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]$ can be expressed as the union of multiple intervals: $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]=\bigcup\limits_{k=1}^{j}(b_{k}-w_{t}^{\prime}\theta_{0},b_{k+1}-w_{t}^{\prime}\theta_{0}]$, the above lower bound can be also written as
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $\geq\sum_{k=1}^{j}\Pr(b_{k}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{k+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$
$=\sum_{k=1}^{j}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
for any $j\leq J$. Fixing the constant $a$, there exists a maximum choice $j^{*}_{a}$, defined as $j^{*}_{a}:=\max\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$. Then, the largest lower bound is the sum of choice probabilities up to the maximum choice $j^{*}_{a}$:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{k=1}^{j^{*}_{a}}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j^{*}_{a}+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
An equivalent expression for this lower bound is given as:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
as the indicator function is zero $\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}=0$ for any choice $j>j^{*}_{a}$.
2. Upper bound. Similarly, we can derive the upper bound for $F_{v_{t}\mid w_{st}}(a\mid w_{st})$ using the sum of conditional choice probabilities, given as follows:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=1-\Pr(v_{t}>a\mid w_{st})$
$\leq 1-\sum_{j=1}^{J}\Pr(b_{j}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}$
$\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}.$
Then, the upper bound for the distribution $F_{v_{t}\mid z_{st}}(a\mid z_{st})$ is derived as follows by taking expectation over covariate $X_{ist}$ given $z_{st}$:
$F_{v_{t}\mid z_{st}}(a\mid z_{st})\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j}-z_{t}^{\prime}\beta_{0}-X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
Given the established lower and bounds on the distribution of $v_{t}\mid z_{st}$, an identified set for $\theta_{0}$ is characterized in the following proposition.

### v3

@@ 2.5 Point Identification
Proposition 1 characterizes the sharp identified set for $\theta_{0}$ by only imposing Assumption 1. This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$ under support conditions on the exogenous covariate $Z_{it}$. We focus on the scenario where the endogenous covariate $X_{it}$ is discrete $X_{it}\in\mathcal{X}\equiv\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ and there are only two periods $T=2$.
@@
$\mathcal{Z}_{1}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\mathbb{P}(Y_{i1}=0,X_{i1}=x\mid z)<\mathbb{P}(Y_{i2}=1,X_{i2}=x\mid z)\Big{\}},$
$\mathcal{Z}_{2}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\mathbb{P}(Y_{i1}=1,X_{i1}=x\mid z)<\mathbb{P}(Y_{i2}=0,X_{i2}=x\mid z)\Big{\}}.$
@@
$(z_{1},z_{2})\in\mathcal{Z}_{3}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}_{4}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}.$
@@
Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}_{3}^{j}\cap\mathcal{Z}_{2}\neq\emptyset$ or $\mathcal{Z}_{4}^{j}\cap\mathcal{Z}_{1}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
When the endogenous variable $X_{it}$ is a scalar, e.g., the lagged dependent variable $X_{it}=Y_{i,t-1}$, then the definition of the two sets $\mathcal{Z}_{3}^{j},\mathcal{Z}_{4}^{j}$ can be simplified as there existing $x_{1}\neq x_{2}$ such that the corresponding choice probability is observed. Besides the sign of $\gamma_{0}$, the identification results can also bound the value of $\gamma_{0}$ from variation in the exogenous covariates.
When $X_{it}$ is multi-dimensional such as including two lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2})$ with $\gamma_{0}=(\gamma_{0}^{1},\gamma_{0}^{2})$, then $\gamma_{0}^{1}$ is identified when the required choice probability in the two sets $\mathcal{Z}_{3}^{1},\mathcal{Z}_{4}^{1}$ are observed for $(Y_{i,1},Y_{i,0})=(1,1),(Y_{i,2},Y_{i,1})=(0,1)$ or $(Y_{i,1},Y_{i,0})=(0,0),(Y_{i,2},Y_{i,1})=(1,0)$. We provide general sufficient conditions to identify the sign of $\gamma_{0}$, which may be stronger than necessary and can be relaxed in certain scenarios. For example, when we know that $\gamma_{0}^{1}+\gamma_{0}^{2}>0$ while $\gamma_{0}^{1}<0$, we can infer that $\gamma_{0}^{2}>0$ without requiring additional assumptions on the two sets $\mathcal{Z}_{3}^{2},\mathcal{Z}_{4}^{2}$.
3 Generalization
3.1 General Identification Strategy
The key idea underlying our identification strategy generalizes further beyond the binary choice model. In this section, we study a general semiparametric model that can accommodate a wide range of panel data models. This general model allows for various types of dependent variables, including ordered, multinomial, and censored outcomes. Moreover, we explore more flexible specifications for the dependent variable, allowing for multidimensional fixed effects and time-varying errors, as well as nonseparable interactions between them.
$Y_{it}$ $=G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right),$ (18)
where $Y_{it}\in\mathcal{Y}$ can be either a discrete or continuous variable, $\alpha_{i}$ is the individual fixed effect of arbitrary dimension, $\epsilon_{it}$ is the time-varying error of arbitrary dimension, $W_{it}$ is a vector of observable covariates, $\theta_{0}\in\mathcal{R}^{d}$ is a conformable vector of parameter, and the function $G$ is allowed to be unknown, nonseparable but assumed to be weakly monotone in the the parametric index:
Assumption 3 (Monotonicity).
Note that, by setting $\alpha_{i},\epsilon_{it}$ to be scalar-valued, and $G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\mathbf{\mathbbm{1}}\left\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\}$, we obtain the binary choice model in Section 2, where $G$ is by construction weakly increasing in $W_{it}^{{}^{\prime}}\theta_{0}$. Beyond binary choice model, since the dependent variable $Y_{it}$ is not restricted, we show that this general model can accommodate a wide range of panel models with various types of dependent variables.
Example 3 (Ordered Response Model).
Consider that the dependent variable $Y_{it}$ can take $J$ possible ordered values: $Y_{it}\in\{y_{1},...,y_{J}\}$ with $y_{j}<y_{j+1}$, and it is generated as follows:
$Y_{it}^{*}$ $=W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},$
@@ 3.1 General Identification Strategy
where the threshold parameters satisfy $b_{1}=-\infty,b_{J+1}=\infty$, and $b_{j}\leq b_{j+1}$.
Then the function $G$ is given as
$G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\sum_{j=1}^{J}y_{j}\mathbf{\mathbbm{1}}\left\{b_{j}<W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\leq b_{j+1}\right\}$
and it is weakly monotone in $W_{it}^{{}^{\prime}}\theta_{0}$ since $y_{j}<y_{j+1}$ for any $11\leq j\leq J-$ 1.
Example 4 (Multinomial Choice Model).
Consider that the dependent variable $Y_{it}$ takes $J$ possible unordered choices: $Y_{it}\in\mathcal{J}=\{0,1,...,J\}$. The latent utility $u_{ijt}$ for each option $j$ and the dependent variable $Y_{it}$ are generated as follows:
$u_{ijt}$ $=W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt},$
$Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt}.$
Although the $J$ choices are unordered and cannot be directly compared, we can construct a new variable $\tilde{Y}_{it}^{K}=\mathbf{\mathbbm{1}}\{Y_{it}\in K\}$ for any subset $K\subset\mathcal{J}$, representing individuals’ choice belonging to the set $K$. It is equivalent to the event that there exists one choice from the set $K$ yielding higher utility than other choices:
$\tilde{Y}_{it}^{K}$ $=\mathbf{\mathbbm{1}}\{W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \forall j\in K,k\in\mathcal{J}\backslash K\}$
$:=G\left((W_{ijt}-W_{ikt})^{\prime}\theta_{0},\alpha_{ij}-\alpha_{ik},\epsilon_{ijt}-\epsilon_{ikt},\ \forall j\in K,k\in\mathcal{J}\backslash K\right).$
Given this new variable $\tilde{Y}_{it}^{K}$, the function $G$ is weakly monotone in $(W_{ijt}-W_{ikt})^{\prime}\theta_{0}$ for $j\in K,k\in\mathcal{J}\backslash K$ and for any set $K\subset\mathcal{J}$, and we have a similar monotonicity structure.
Example 5 (Censored Outcome Model).
Consider the dependent variable $Y_{it}$ is censored at 0, given as follows:
$Y^{*}_{it}$ $=W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
and the function $G(W_{it}^{{}^{\prime}}\theta_{0},\alpha_{i},\epsilon_{it})=\max\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},0\}$ clearly satisfies the monotonicity in $W_{it}^{{}^{\prime}}\theta_{0}$.
We have shown that the semiparametric model with monotonicity can accommodate various types of panel data models. Now we describe our identification strategy for this general model. As before, we decompose the covariate $W_{it}$, and correspondingly $\theta_{0}$, into two components, $W_{it}=\left(Z_{it}^{{}^{\prime}},X_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $\theta_{0}=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$so that
where $Z_{it}$ denotes exogenous covariates while $X_{it}$ denotes endogenous covariates, with the precise definition of exogeneity encoded by the partial stationarity in Assumption 1. We show how partial stationarity can be exploited in conjunction with weak monotonicity (Assumption 3) to obtain identifying restrictions in the presence of endogeneity.
Let $\mathcal{Y}$ denote the support of $Y_{it}$. For any $c\in\mathcal{R}$ and $y\in\mathcal{Y}$, conditional on $W_{ist}=w_{st}$, we consider the event that
$Y_{it}\leq y,\quad\quad Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c,$
implying that
$\mathbf{\mathbbm{1}}\{Y_{it}\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\}$ $=\mathbf{\mathbbm{1}}\left\{G\left(Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},\,\alpha_{i},\,\epsilon_{it}\right)\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}$
$\leq\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\},$
where the inequality holds by the monotonicity assumption on the function $G$.
Symmetrically, we can provide an upper bound for $\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}$ by looking at the following event:
$\mathbf{\mathbbm{1}}\{Y_{it}>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\}$ $=\mathbf{\mathbbm{1}}\left\{G\left(Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},\,\alpha_{i},\,\epsilon_{it}\right)>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$
$\leq\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)>y\}$
$=1-\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}.$
which is equivalent to
$\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}\leq 1-\mathbf{\mathbbm{1}}\{Y_{it}>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\}.$
The partial stationarity assumption in 1 implies that
$\mathbb{P}\left(Y_{it}\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\mid Z_{i}=z\right)$
$=$ $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\mid Z_{i}=z\right)$
$=$ $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{is}\right)\leq y\mid Z_{i}=z\right)$
$\leq$ $1-\mathbb{P}\left(Y_{is}>y,\ Z_{is}^{{}^{\prime}}\beta_{0}+X_{is}^{{}^{\prime}}\gamma_{0}\leq c\mid Z_{i}=z\right).$
The key difference of the above identification strategy in (LABEL:eq:Bounds_z) relative to the corresponding identifying restrictions in previous sections lies in that the “middle term” in (LABEL:eq:Bounds_z) is no longer the conditional CDF of $\alpha_{i}+\epsilon_{it}$, but a conditional distribution of $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{is}\right)\leq y\mid Z_{i}=z\right)$, with the latter representation not dependent on scalar-additivity of fixed effects and time-varying errors.
Proposition 5 (Identified Set).
Under Assumptions 1 and 3, $\theta_{0}\in\Theta_{I,gen}$, where the identified set $\Theta_{I,gen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$\max_{t=1,...,T}\mathbb{P}\left(Y_{it}\leq y,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\mid Z_{i}=z\right)\leq 1-\max_{s=1,...,T}\mathbb{P}\left(Y_{is}>y,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\leq c\mid Z_{i}=z\right),$ (19)
for any $c\in\mathcal{R},$ $y\in\mathcal{Y}$, and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
Note that in the binary choice setting of Section 2, it suffices to set $y=0$ in (19), which then coincides with the identifying results in Proposition 1. More generally, we show that our identification results in Proposition 5 can be adapted to the ordered response model in Section 3.2, the multinomial choice model in Section 3.3, and the censored outcome model in Section 3.4.
The results in Proposition 5 generally hold regardless of whether the dependent variable and the endogenous covariate are discrete or continuous. The next proposition shows that additional discreteness in either the dependent variable or endogenous covariates can further simplify and reduce the number of the identifying conditions in (19).
Proposition 6 (Discreteness).
When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for any $t$, then $\Theta_{I,gen}=\Theta_{I,gen}^{disc_{x}},$ where $\Theta_{I,gen}^{disc_{x}}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}$ that satisfy condition (19) for any $c\in\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\}$.
Moreover, when $Y_{it}\in\left\{\overline{y}_{1},...,\overline{y}_{K}\right\}$ with $\overline{y}_{j}\leq\overline{y}_{j+1}$ for any $t$, then $\Theta_{I,gen}=\Theta_{I,gen}^{disc_{y}},$ where $\Theta_{I,gen}^{disc_{y}}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}$ that satisfy condition (19) for any $y\in\left\{\overline{y}_{1},...,\overline{y}_{K-1}\right\}$.
Proposition 6 shows that for any discrete choice models with discrete endogenous variables, such as dynamic binary, ordered, and multinomial choice models, the identified set $\Theta_{I,gen}$ is characterized by a finite number of moment inequalities. Sections 3.2-3.4 establish general identification results for panel models with various types of dependent variables and endogeneity, and also explore both the static model without endogeneity and the dynamic model with lagged dependent variables.
3.2 Ordered Response Model
Consider that the outcome variable $Y_{it}$ takes $J$ ordered values: $Y_{it}\in\{y_{1},..,y_{J}\}$ with $y_{j}\leq y_{j+1}$. Examples of such ordered outcomes include various income categories, health outcomes, or levels of educational attainment. We study the following panel ordered choice model:
$Y_{it}^{*}$ $=W_{it}^{\prime}\theta_{0}+\alpha_{i}+\epsilon_{it},$ (20)
$Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\},$
@@ 3.2 Ordered Response Model
where $Y_{it}^{*}$ denotes the latent dependent variable, and $Y_{it}$ denotes the ordered outcome which takes value $y_{j}$ when $Y_{it}^{*}\in(b_{j},b_{j+1}]$. The threshold parameters satisfy $b_{1}=-\infty,b_{J+1}=+\infty$, and the remaining threshold parameters $b_{j}$ (where $b_{j}\leq b_{j+1}$) can be either known or unknown for $12\leq j\leq J-$ 1. The binary choice model in (1) is nested with $J=2$ and $b_{2}=0$.
For this ordered choice model, we observe the conditional probability of each choice $j$:
$\mathbb{P}\left(Y_{it}\leq y_{j}\mid w\right)=\mathbb{P}\left(\alpha_{i}+\epsilon_{it}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w\right),$
and this choice probability is monotone with respect to $b_{j+1}-w_{t}^{\prime}\theta_{0}$. Compared to the binary choice model, here we can exploit variations in the conditional probability across all possible choices to provide more informative results for $\theta_{0}$.
Following the same identification strategy in Sections 2 and 3, we can bound the conditional probability of $\mathbb{P}\left(\alpha_{i}+\epsilon_{it}\leq c\mid w\right)$ below as follows: for any constant $c$,
$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\geq\mathbb{P}\left(Y_{it}\leq y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right).$ (21)
Since the above inequality holds for any choice $j$, we can take the largest lower bound over all possible choices that satisfy $b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c$. We define $\overline{j}_{c}$ as $\overline{j}_{c}:=\max\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\}$, and condition (21) leads to:
$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$ $\geq\sum_{k=1}^{\overline{j}_{c}}\mathbb{P}\left(Y_{it}=y_{k},\ b_{\overline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right)$
$=\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right).$
where the last inequality holds since $\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\}=0$ for any choice $j>\overline{j}_{c}$. Furthermore, the final expression offers the advantage of avoiding the need to search for the maximum choice $\overline{j}_{c}$ for each constant $c$.
Similarly, we can derive an upper bound for the probability $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$:
$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\leq 1-\mathbb{P}\left(Y_{it}>y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right).$
Define $\underline{j}_{c}$ as $\underline{j}_{c}:=\min\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\geq c\}$, then we have
$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$ $\leq 1-\mathbb{P}\left(Y_{it}>b_{\underline{j}_{c}},\ b_{\underline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right)$
$=1-\mathbb{P}\left(Y_{it}\geq b_{\underline{j}_{c}+1},\ b_{\underline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right).$
Furthermore, the above inequality can be equivalently written as
$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\leq 1-\sum_{j=1}^{J}\mathbb{P}\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq c\mid z\right),$
which holds since for $\mathbf{\mathbbm{1}}\{b_{j+1}-w_{t}^{\prime}\theta_{0}<c\}=0$ for any choice $j<\underline{j}_{c}$.
Given the established lower and upper bounds on the conditional probability $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$, we can derive the corresponding bounds for $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid z)$ by taking expectation over the endogenous covariate $X$. Then, the identifying condition for $\theta_{0}$ is characterized by the restriction that the bounds over different periods must have nonempty intersections, as presented in the following proposition.